# Training-Free Stylized Abstraction

### v1

@@ 1 Introduction
Image-to-image style translation deng2022stytr2 ; sohn2023styledrop ; wang2023stylediffusion ; jiang2024diffartist ; jing2019neural ; xing2024csgo ; chen2021artistic is a well-studied area that traces its origins to GAN-based approaches, such as neural style transfer gatys2015neural and CycleGAN zhu2017unpaired , and has since evolved to include both diffusion-based training and training-free methods rout2024semantic ; wang2024taming ; le2022styleid ; mo2024freecontrol ; zhang2023adding ; zhao2023uni ; deng2022stytr2 ; liu2023name ; chen2024artadapter . These techniques typically focus on overlaying a specific style onto an input image while preserving the subject’s identity. Common examples include transforming portraits into sketches, cartoons, or artwork in the style of artists like Van Gogh. Importantly, the resulting stylized images often retain structural consistency with the original content.
@@
To address the limitations of existing stylization methods, we present a training-free framework for stylized abstraction that generalizes beyond celebrity likenesses to everyday identities and supports a wide range of abstract styles: including LEGO, South Park, Simpson, Matrushka, Barbie, Knitted Doll, Action Figure, etc. Our approach requires neither subject-specific fine-tuning nor dataset-level adaptation. Instead, we introduce a novel inference-time scaling strategy for vision-language models (VLLMs) that distills core semantic traits critical for identity preservation and aligns them with user-driven stylistic prompts. Central to our pipeline is a multi-turn generation loop, where missing or distorted identity cues, identified via VLLM feedback, are progressively reintegrated to enhance fidelity across iterations. To recover subject structure under extreme abstraction, we extend RF-Inversion rout2024semantic with a cross-domain latent inversion scheme, treating stylized images as source latents and photorealistic representations as structural targets. Leveraging rectified flow-guided updates and style-aware temporal scheduling, our method preserves stylistic fidelity while selectively restoring identity-consistent structure in a controllable and interpretable fashion. To evaluate abstraction beyond pixel-level similarity, we introduce StyleBench, a GPT-assisted, human-aligned protocol for benchmarking stylized abstraction. We further report quantitative performance using KID and CLIPScore, supported by a user preference study. Our framework sets a new state-of-the-art in abstraction quality: fully training-free, identity-consistent, and broadly generalizable across styles and subjects.
@@ 2 Related Work
Identity-Preserving Style Transfer. Identity-preserving or subject-driven style transfer zhang2024ssr ; raj2023dreambooth3d ; chen2023subject ; miao2024subject ; dong2022dreamartist ; han2023highly ; voynov2023p+ ; alaluf2023neural ; kumari2023multi ; liu2023cones ; han2023svdiff ; ryu2023low ; avrahami2023break ; chen2023disenbooth ; cai2024decoupled ; he2025data is a closely related line of work to stylized abstraction, where the goal is to synthesize stylized images while retaining subject identity. Approaches in this domain broadly fall into three categories. The first category includes fine-tuning-based methods, such as Textual Inversion gal2022image and DreamBooth ruiz2023dreambooth , which adapt the generative model to the target subject using multiple reference images. While effective for object-centric domains, these methods often struggle to faithfully preserve identity in human subjects and require several input images for fine-tuning. The second category comprises encoder-based methods that learn feature adaptation modules to modulate the base model without explicit fine-tuning. Notable examples include IP-Adapter ye2023ip and BLIP-Diffusion li2023blip , which leverage pretrained encoders to align content and style representations. The third category focuses on single-image personalization, including DreamTuner hua2023dreamtuner and CSGO xing2024csgo , or entirely training-free techniques such as RF-Inversion rout2024semantic , RF-Modulation wang2024taming , StyleID le2022styleid , InstantID wang2024instantid , InstantID-Plus wang2024instantstyle , and DiffArtist jiang2024diffartist . These methods often rely on CLIP-guided optimization or feature injection to steer generation toward the desired style. However, such methods typically prioritize structural fidelity or stylization strength in isolation. In high stylization scenarios, they may struggle to abstract and reinterpret content meaningfully, as their architectures tend to enforce either identity preservation or stylistic consistency without deeper semantic understanding. Bridging this gap requires novel approaches that reason over semantic correspondences between style and identity, rather than relying solely on pixel or feature-space alignment.
@@ 3.2 Cross-Domain Latent Reversal with Rectified Flows
At this stage, we obtain a highly stylized representation of the subject with faithfully preserved stylistic elements, $y_{s}$, but the generation now requires structural guidance from the original image to recover key identity-aligned geometry. However, unlike prior inversion-based pipelines rout2024semantic that operate on realistic or noisy inputs, our setting begins from an already abstracted, stylized image $y_{s}$. This shift introduces a novel challenge: how to reconstruct a semantically grounded latent representation from a highly altered input while flexibly recovering structural details based on style demands.
To address this, we propose a two-stage framework for cross-domain latent reversal, which extends rectified flow methods rout2024semantic into the abstract stylization space. The key is to treat the stylized abstraction not as a degraded variant of a photo, but as a valid starting point in an altered visual domain: one that must be softly regularized back into a structured latent aligned with the true subject identity.
@@ 4 Experiment
Baselines There is no direct baseline for stylized abstraction, as it is a relatively new concept in the vision community that goes beyond identity preservation to include semantic and geometric reinterpretation. We compare against the closest related methods across personalization and style transfer. These include fine-tuning-based approaches such as Textual Inversion gal2022image and DreamBooth ruiz2023dreambooth , encoder-based CSGO xing2024csgo , and training-free, zero-shot methods including StyleID le2022styleid , RF-Inversion rout2024semantic , RF-Modulation wang2024taming , DiffArtist jiang2024diffartist , InstantID wang2024instantid , and InstantID-Plus wang2024instantstyle .
@@ 5 Results and Analysis
RF-Inversion rout2024semantic 0.166 0.1902 1.5 2
RB-Modulation wang2024taming 0.035 0.2069 0.5 0.5
@@ Appendix B Baselines Setup
RF-Inversion [38]. For RF-Inversion experiments, we use the FluxPipeline with pretrained weights from FLUX.1-dev, applying a fixed set of 10 stylized prompts (e.g., Knitted Doll, Lego, Anime, Pixar, Barbie Doll, etc.). Each input image is first inverted into the latent space using 28 inversion steps with a gamma value of 0.5. Stylized outputs are then generated from the inverted latents using the corresponding prompt, with a diffusion range restricted to the first 25% of the process (start timestep=0.0, stop timestep=0.25) across 50 inference steps. We use eta=0.9 for sampling noise and fix the random seed for deterministic results.

### v2

@@ 1 Introduction
Image-to-image style translation deng2022stytr2 ; sohn2023styledrop ; wang2023stylediffusion ; jiang2024diffartist ; jing2019neural ; xing2024csgo ; chen2021artistic is a well-studied area that traces its origins to GAN-based approaches, such as neural style transfer gatys2015neural and CycleGAN zhu2017unpaired , and has since evolved to include both diffusion-based training and training-free methods rout2025semantic ; rout2025rbmodulation ; le2022styleid ; mo2024freecontrol ; zhang2023adding ; zhao2023uni ; deng2022stytr2 ; liu2023name ; chen2024artadapter . These techniques typically focus on overlaying a specific style onto an input image while preserving the subject’s identity. Common examples include transforming portraits into sketches, cartoons, or artwork in the style of artists like Van Gogh. Importantly, the resulting stylized images often retain structural consistency with the original content.
@@
To address the limitations of existing stylization methods, we present a training-free framework for stylized abstraction that generalizes beyond celebrity likenesses to everyday identities and supports a wide range of abstract styles: including LEGO, South Park, Simpson, Matrushka, Barbie, Knitted Doll, Action Figure, etc. Our approach requires neither subject-specific fine-tuning nor dataset-level adaptation. Instead, we introduce a novel inference-time scaling strategy for vision-language models (VLLMs) that distills core semantic traits critical for identity preservation and aligns them with user-driven stylistic prompts. Central to our pipeline is a multi-turn generation loop, where missing or distorted identity cues, identified via VLLM feedback, are progressively reintegrated to enhance fidelity across iterations. To recover subject structure under extreme abstraction, we extend RF-Inversion rout2025semantic with a cross-domain latent inversion scheme, treating stylized images as source latents and photorealistic representations as structural targets. Leveraging rectified flow-guided updates and style-aware temporal scheduling, our method preserves stylistic fidelity while selectively restoring identity-consistent structure in a controllable and interpretable fashion. To evaluate abstraction beyond pixel-level similarity, we introduce StyleBench, a GPT-assisted, human-aligned protocol for benchmarking stylized abstraction. We further report quantitative performance using KID and CLIPScore, supported by a user preference study. Our framework sets a new state-of-the-art in abstraction quality: fully training-free, identity-consistent, and broadly generalizable across styles and subjects.
@@ 2 Related Work
Identity-Preserving Style Transfer. Identity-preserving or subject-driven style transfer zhang2024ssr ; raj2023dreambooth3d ; chen2023subject ; miao2024subject ; dong2022dreamartist ; han2023highly ; voynov2023p+ ; alaluf2023neural ; kumari2023multi ; liu2023cones ; han2023svdiff ; ryu2023low ; avrahami2023break ; chen2023disenbooth ; cai2024decoupled ; he2025data is a closely related line of work to stylized abstraction, where the goal is to synthesize stylized images while retaining subject identity. Approaches in this domain broadly fall into three categories. The first category includes fine-tuning-based methods, such as Textual Inversion gal2022image and DreamBooth ruiz2023dreambooth , which adapt the generative model to the target subject using multiple reference images. While effective for object-centric domains, these methods often struggle to faithfully preserve identity in human subjects and require several input images for fine-tuning. The second category comprises encoder-based methods that learn feature adaptation modules to modulate the base model without explicit fine-tuning. Notable examples include IP-Adapter ye2023ip and BLIP-Diffusion li2023blip , which leverage pretrained encoders to align content and style representations. The third category focuses on single-image personalization, including DreamTuner hua2023dreamtuner and CSGO xing2024csgo , or entirely training-free techniques such as RF-Inversion rout2025semantic , RB-Modulation rout2025rbmodulation , StyleID le2022styleid , InstantID wang2024instantid , InstantID-Plus wang2024instantstyle , and DiffArtist jiang2024diffartist . These methods often rely on CLIP-guided optimization or feature injection to steer generation toward the desired style. However, such methods typically prioritize structural fidelity or stylization strength in isolation. In high stylization scenarios, they may struggle to abstract and reinterpret content meaningfully, as their architectures tend to enforce either identity preservation or stylistic consistency without deeper semantic understanding. Bridging this gap requires novel approaches that reason over semantic correspondences between style and identity, rather than relying solely on pixel or feature-space alignment.
@@ 3.2 Cross-Domain Latent Reversal with Rectified Flows
At this stage, we obtain a highly stylized representation of the subject with faithfully preserved stylistic elements, $y_{s}$, but the generation now requires structural guidance from the original image to recover key identity-aligned geometry. However, unlike prior inversion-based pipelines rout2025semantic that operate on realistic or noisy inputs, our setting begins from an already abstracted, stylized image $y_{s}$. This shift introduces a novel challenge: how to reconstruct a semantically grounded latent representation from a highly altered input while flexibly recovering structural details based on style demands.
To address this, we propose a two-stage framework for cross-domain latent reversal, which extends rectified flow methods rout2025semantic into the abstract stylization space. The key is to treat the stylized abstraction not as a degraded variant of a photo, but as a valid starting point in an altered visual domain: one that must be softly regularized back into a structured latent aligned with the true subject identity.
@@ 4 Experiment
Baselines There is no direct baseline for stylized abstraction, as it is a relatively new concept in the vision community that goes beyond identity preservation to include semantic and geometric reinterpretation. We compare against the closest related methods across personalization and style transfer. These include fine-tuning-based approaches such as Textual Inversion gal2022image and DreamBooth ruiz2023dreambooth , encoder-based CSGO xing2024csgo , and training-free, zero-shot methods including StyleID le2022styleid , RF-Inversion rout2025semantic , RB-Modulation rout2025rbmodulation , DiffArtist jiang2024diffartist , InstantID wang2024instantid , and InstantID-Plus wang2024instantstyle .
@@ 5 Results and Analysis
RF-Inversion rout2025semantic 0.166 0.1902 1.5 2
RB-Modulation rout2025rbmodulation 0.035 0.2069 0.5 0.5
@@ Appendix B Baselines Setup
RF-Inversion [rout2024semantic]. For RF-Inversion experiments, we use the FluxPipeline with pretrained weights from FLUX.1-dev, applying a fixed set of 10 stylized prompts (e.g., Knitted Doll, Lego, Anime, Pixar, Barbie Doll, etc.). Each input image is first inverted into the latent space using 28 inversion steps with a gamma value of 0.5. Stylized outputs are then generated from the inverted latents using the corresponding prompt, with a diffusion range restricted to the first 25% of the process (start timestep=0.0, stop timestep=0.25) across 50 inference steps. We use eta=0.9 for sampling noise and fix the random seed for deterministic results.